A circular coil of radius 5x10-2m and with 40 turns is carrying a current of 0.25 A. determine the magnetic field of the circular coil at the center.


1
Expert's answer
2022-03-15T10:34:35-0400

The magnetic field at the center of circular coil

"B=\frac{\mu_0 N I}{2\pi r}"

"B=\frac{4\pi*10^{-7}*40*0.25}{2\pi*5*10^{-2}}=4*10^{-5}\:\rm T"